A 1.30 kg squirrel runs east across the ground for 6.00 meters. She climbs 0.600 meters before running west on a limb for 3.00 meters. She jumps up .900 meters onto another branch, travels west 4.25 meters. She stops and eats a nut. Calculate net work. What is the squirrel’s gravitational potential energy?
1
Expert's answer
2020-11-23T10:29:06-0500

The net work against gravity is given by:


"A = mgh"

where "m = 1.3kg" is the mass of the squirrel, "g = 9.81m\/s^2" is the gravitational acceleration and "h" is the final height of the squirrel above the ground:


"h = 0.6m + 0.9m = 1.5m"

Thus, obtain:


"A = 1.3\cdot 9.81\cdot 1.5 \approx 19.13J"

The final potential energy of the squirrel is the same:


"U = A = 19.13J"

Answer. 19.13J, 19.13J.